OpenStudy (anonymous):

12. Which equation is an identity? (1point) A.8-(5x + 2)= 5x - 6 B.7z + 10 - z = 8z - 2(z-5) C.8m - 4 = 5m + 8 - m D.6y + 5 = 6y - 5 13. Which equation has no solution? (1point) A. 7v + 2 = 8v - 3 B. 3x - 5 = 3x + 8 - x C. 4y + 5 = 4y - 6 D. 7z + 6 = -7z - 5 14. Solve the following equation (1point) 5(z - 4) - z = 4z - 20 A. 4 B. -4 C. Infinitely many solutions D. No solution 15. Solve the following equation (1point) 4 + 2y = 9 + 2y A. 4 B. -4 C. Infinitely many solu

OpenStudy (anonymous):

@didithedad

OpenStudy (anonymous):

Sorry the reply took so long. An identity is an equation which is true for every value of the variable is called an identity equation. To make it shorter, I will do 1 example. B says that 7z + 10 - z = 8z - 2(z-5) Well 7z - z would be 6z + 10. the other equation is 8z - 2z-10 simplified it is 6z-10 6z-10 and 6z + 10 are not the same, so it is not b. The answer is C.

OpenStudy (anonymous):

Thanks dude

OpenStudy (anonymous):

Skipping down to 14. 5(z - 4) - z = 4z - 20 Simplify 5z - 20 - z = 4z - 20 4z - 20 = 4z - 20 The answer would be infinite. No matter what you put in for z, the answer will always be the same. C.

OpenStudy (anonymous):

#15 4 + 2y = 9 + 2y No solutions.

OpenStudy (anonymous):

13 is z.

OpenStudy (anonymous):

Sorry, i meant C

OpenStudy (anonymous):

So 12= C 13= C 14= C 15= no solutions

OpenStudy (anonymous):

not completely 100% on those answers, but its the best I can do. :)